OpenStudy (anonymous):

A circle has its center at (-1, 2) and a radius of 3 units. What is the equation of the circle?

OpenStudy (anonymous):

(x + 1)2 + (y - 2)2 = 9 (x - 1)2 + (y + 2)2 = 3 (x + 1)2 + (y - 2)2 = 3 (x + 1)2 + (y + 2)2 = 9

OpenStudy (dumbcow):

plug it into the formula --> (x-h)^2 + (y-k)^2 = r^2

OpenStudy (dumbcow):

(h,k) is center r is radius

OpenStudy (anonymous):

take a guess... which one?

OpenStudy (anonymous):

use what dumbcow says...

OpenStudy (anonymous):

the last one?

OpenStudy (anonymous):

guess again...

OpenStudy (anonymous):

first one